# Elongated vortex quantum droplets in binary Bose-Einstein condensates

### v1

@@ Abstract
Stability of elongated (“slender”) quantum droplets (QDs) with embedded unitary and multiple vorticity is a problem that was not solved previously. In this work, we propose a solution which relies upon the use of the spatial modulation of the inter-species scattering length in the binary Bose-Einstein condensate, in the form of a two-dimensional axisymmetric Gaussian, shaped by means of the optical Feshbach resonance. The corresponding effective nonlinear trapping potential supports completely stable elongated QDs with vorticity $S=0$ and partly stable families of elongated QDs with $S=1,2,3,4$ (other nonlinear systems do not maintain stability of vortex droplets with $\geq 2$). We systematically analyze effects of the amplitude and width of the Gaussian modulation, as well as the total number of atoms, on the shape and stability of the QDs, some effects being explained analytically. Collisions between identical QDs with $S=1$ moving in opposite directions along the central axis leads to their merger into still more elongated breathing QDs with the same vorticity, while collisions between QDs with $S=\pm 1$ are quasi-elastic. Moving modulation profiles are able to adiabatically rotate the trapped elongated QDs. Application of a torque to the vector QD sets in the gyroscopic regime of robust precession, which realizes a macroscopic spin-orbit-coupling effect.
@@
Introductions and the model. Vortices with an elongated three-dimensional (3D) structure have drawn much interest in astrophysics McWilliams1994a ; astro and geophysics McWilliams1994b , plasmas plasma1 ; plasma2 ; plasma3 , superconductors Dolan1989 ; Odobesko2020 ; Hung2012 ; Olszewski2018 , optics Freund1997 ; Sharma2011 , and Bose-Einstein condensates (BECs) Stockhofe2011 ; Liu2013 ; McEndoo2010 ; OKtel2004 ; Engels2002 ; Tikhonenkov2008 ; Tan2023 . A powerful platform for the realization of various dynamical states in binary BEC is provided by the quantum droplets (QDs), which are stabilized by the beyond-mean-field (Lee-Huang-Yang, LHY) effect of quantum fluctuations LHY ; Petrov2015 ; Petrov2016 ; Luca1 ; Jrgensen2018 ; Cabrera2018 ; Ferrier-Barbut2018 ; Cheiney2018 ; Inguscio2018 ; Luca2 ; Yin2021 ; dlw2021 ; Luo2021 ; Cui2025 . The isotropy of contact interactions in the binary BEC makes it possible to predict stable 2D and 3D axisymmetric QDs with embedded vorticity and more complex vortex patterns Kartashov2018 ; QDvortex2D ; Tengstrand2019 ; Chen2024 ; Li2024FOPview .
Dynamics of BECs of another type, which are composed of atoms carrying magnetic moments polarized by an external magnetic field, is dominated by the long-range dipole-dipole interactions (DDIs) Stuhler2005 ; Lahaye2009 . In the combination with the beyond-mean-field effects, the naturally anisotropic DDIs support single-component QDs which feature strongly anisotropic density distributions Ferrier-Barbut2016 ; Schmitt2016 ; Santos ; Wang2020 ; Bottcher2021 ; Yin2022 . Stable 3D anisotropic vortex QDs in dipolar BEC, in which the vorticity axis is perpendicular to the direction of the dipolar polarization, have been predicted too Li2024PRL ; Li2024FOP , while the search for elongated QDs that maintain the rotational symmetry about the vortex axis aligned with the polarization produces unstable modes Cidrim2018 .
While, as said above, 3D axisymmetric QDs were predicted in the non-dipolar binary BEC with the isotropic contact interactions Kartashov2018 , they are stable only in the limit case of a very large size (which corresponds to an extremely large number of atoms, $\sim 5\cdot 10^{5}$ Kartashov2018 ), making the creation of such stable vortex QDs a challenging task for the experiment. More compact elongated QDs with embedded vorticity, that might be a more appropriate target for experiments, were predicted too as stationary states; however, as mentioned above too, they are unstable in the dipolar and non-dipolar BEC alike Kartashov2018 ; Cidrim2018 . The objective of the present work is to predict stable vortex QDs with a “slender” (elongated) shape in the binary condensate with the contact interactions, using an effective nonlinear potential RMP (alias pseudopotential pseudo ), that may be induced by a spatially modulated interaction strength. In turn, the modulation can be imposed by means of the spatially inhomogeneous Feshbach resonance. In its usual spatially uniform implementation, the Feshbach resonance is induced by dc magnetic field which alters the s-wave scattering length $a_{s}$ of the inter-atomic interaction. This effect is widely employed as a standard tool for manipulating the intrinsic nonlinearity in BEC Tiesinga1993 ; Moerdijk1995 ; Courteille1998 ; Bertozzi1998 ; Inouye1998 ; Timmermans1999 ; Duine2004 ; kohler2006 ; Jones2006 ; Randy ; Chin2010 . A similar method can be applied in the form of the optical Feshbach resonance (OFR) Fisher1989 ; Fedichev1996 ; Caputo ; Saito2003 ; Qi2011 ; Chien2012 . In particular, OFR can be readily induced by spatially non-uniform optical fields with inhomogeneous intensity $I\left(x,y,z\right)$, which can be used to shape effective nonlinear potentials through the local OFR relation,
@@
where $a_{\mathrm{opt}}$ is the OFR-induced change of the scattering length, $\left(x,y,z\right)$ are spatial coordinates, and sign minus implies that the resonant laser illumination induces the inter-atomic attraction force Theis2004 ; Bauer2009 ; Yamazaki2010 ; RMP ; Yan2013 .
@@
where $L$ is the transverse radius of the beam, and $-a^{(0)}<0$ is the amplitude of the OFR-induced modulation of the local scattering length. The objective is to use the axisymmetric profile (2) to stabilize vortex solitons with the angular momentum directed along the $z$ axis and a “slender” (elongated) shape. The setting will be also used to study more sophisticated dynamical regimes, such as gyroscopic motion (precession) of the vortex soliton.
The dynamics of the 3D binary condensate is governed by the system of Gross-Pitaevskii (GP) equations with the LHY corrections, written with wave functions $\psi_{1,2}$ of the two components:
$i\hbar\frac{\partial\psi_{j}}{\partial t}$ $=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi_{j}+(\mathcal{G}_{jj}|\psi_{j}|^{2}+\mathcal{G}_{j(3-j)}|\psi_{3-j}|^{2})\psi_{j}$
$+\Gamma_{j}(|\psi_{j}|^{2}+|\psi_{3-j}|^{2})^{3/2}\psi_{j},\ \ \ j=1,2,$ (3)
where $\hbar$ and $m$ are the Planck’s constant and atomic mass, while $\mathcal{G}_{jj}=4\pi\hbar^{2}a_{jj}/{m}$ ($j=1,2$) and $\mathcal{G}_{12}=4\pi\hbar^{2}a_{12}/{m}$ are the strengths of the intra- and inter-component mean-field interactions, respectively. Here , $a_{jj}$ $(j=1,2)$ and $a_{12}$ are the corresponding $s$-wave scattering lengths of atomic collisions, and the strengths of the LHY corrections are $\Gamma_{j}=4m^{3/2}\mathcal{G}_{jj}^{5/2}/3\pi^{2}\hbar^{3}$ Petrov2015 ; Fischer2006 ; Inguscio2018 .
We consider the setup with the inter-component scattering length engineered by means of OFR as $a_{12}=a_{\mathrm{bg}}+a_{\mathrm{opt}}$, where $a_{\mathrm{bg}}$ is the background value of the scattering length in the absence of light, which is adjusted by the additional magnetic Feshbach resonance as $a_{\mathrm{bg}}=-\sqrt{a_{11}a_{22}}$, so as to cancel the spatially uniform value of the scattering length for the symmetric system, with $\psi_{1}=\psi_{2}\equiv\psi/\sqrt{2}$, which may replace the GP-LHY system (3) in the case of $a_{11}=a_{22}$, $\Gamma_{1}=\Gamma_{2}\equiv\Gamma$. Thus, the system is replaced by the single GP-LHY equation:
@@
with angular coordinate $\theta$ and real function $\phi\left(\rho,z\right)$ satisfying the stationary equation
@@
The results. To quantify the slenderness of 3D vortex QDs produced by Eq. (4), we define the aspect ratio, $\mathcal{A}=D_{\mathrm{v}}/D_{\mathrm{h}}$, where $D_{\mathrm{v}}$ and $D_{\mathrm{h}}$ are, respectively, vertical and horizontal lengths of the QDs. First, the numerical solution of the stationary version of Eq. (4) for fundamental QDs [with $S=0$ in Eq. (5)] demonstrates, in Figs. 1(a1-a3), the effect of the nonlinear trapping potential on the QD’s shape. In Fig. 1(a1), the aspect ratio is $\mathcal{A}=1$ for the isotropic fundamental QD in the free space ($L=\infty$), and $(a^{(0)},N)=(20a_{0},10^{6})$ ($a_{0}$ is the Bohr radius). As shown in Figs. 1(b1,b2), the aspect ratio increases (i.e., the fundamental QD elongates) with the decrease of radius $L$ of the holding beam in Eq. (2) and increase of $N$. The numerical solution of the time-dependent equation (4) (not shown here in detail) demonstrates that the entire family of the elongated QDs with $S=0$ is stable. Note, in particular, that the $\mu(N)$ dependence plotted in Fig. 1(b2) satisfies the well-known necessary stability condition, $d\mu/dN<0$ (the Vakhitov-Kolokolov criterion Vakhitov1973 ; Berge ).
@@
The independence of the modulation profile (2) on $z$ makes it possible to generate QDs freely moving along the $z$ axis, and thus simulate Eq. (4) for collisions between the QDs moving in opposite directions, initiated by the input, $\psi_{0}\left(\rho,z\right)=\psi(\rho,z-z_{d})\mathrm{e}^{-i\eta z}+\psi(\rho,z+z_{d})\mathrm{e}^{i\eta z}$, where $2z_{d}$ is the initial separation between the QDs, and ±$\pm\eta$ are the opposite kicks which set the droplets in motion. Figures 3(a1-a4) show that the collision between the QDs with identical vorticities $S=1$ is inelastic, resulting in the merger of the vortex QDs into a single oscillatory one (breather). The breathing dynamics is robust and quasi-harmonic, lasting, at least, for $>100\ \mathrm{ms}$, as is illustrated in Fig. 3(b) a sequence of $5$ cycles within the time interval of $50\thicksim 75\ \mathrm{ms}$. The breathing frequency is nearly independent of the collision velocity, suggesting that it is an eigenfrequency of a QD’s intrinsic mode. The collision leads to full destruction of the vortex QDs if the initial kick exceeds a certain critical value, which is $\eta_{\rm{cr}}=1.2\ \upmu\mathrm{m^{-1}}$. Collisions between zero-vorticity QDs produce similar results (not shown here in detail). On the other hand, for the QDs with opposite vorticities, $S=\pm 1$, the simulations demonstrate that the collision is quasi-elastic, leading to temporary fusion into a single droplet, which then splits back into original moving vortex QDs. The latter result is explained by the fact that the opposite vorticities attenuate the coherent interaction between the colliding objects. It is shown in Fig. 3(c) that the holding beam rotating around the $y$-axis in the $\left(xz\right)$-plane at angular velocity $\omega=0.163\ \mathrm{rad/ms}$, which means the replacement of the trapping profile (2) by
@@
Figure 4: (a-c) Typical examples of the vortex droplets and their phases diagrams (at $z=0$) with $(L,a^{(0)},N)=(8\ \upmu\mathrm{m},20a_{0},10^{6})$. (d) The chemical potential $\mu$ of the QDs with $S=0\sim 4$ vs. $N$, with solid and dashed lines representing stable and unstable solutions, respectively.
@@
Conclusion We have proposed to use the OFR, modulating the inter-component atomic $s$-wave scattering length in the binary BEC, to induce the effective 2D nonlinear trapping potential with the Gaussian profile. It supports elongated QDs, including ones with embedded integer vorticity $S$, which are stabilized by the GP-LHY equations. We have systematically explored the effects of the modulation amplitude and width, as well as the total number of atoms, on the shape and stability of the QDs. The fundamental QDs with $S=0$ are completely stable, while families of the vortex QDs with $S=1,2,3,$ and $4$ are partly stable. Collisions of identical elongated vortex QDs with $S=1$ lead to their merger into breathers with the same vorticity, $S=1$, while collisions between the elongated QDs with opposite vorticities, $S=\pm 1$, are quasi-elastic. The moving OFR modulation profile can rotate the elongated QDs in the abiabatic regime. The application of the torque to the trapped elongated vortex QD sets it in a state of stable gyroscopic motion (precession).
As an extension of the analysis, it may be interesting to consider more complex holding potentials, such as anisotropic ones, as well as those combining the nonlinear trap and vortex laser beams.

### v2

@@ Abstract
Stability of elongated (“slender”) quantum droplets (QDs) with embedded unitary and multiple vorticity is a problem that was not solved previously. In this work, we propose a solution which relies upon the use of the spatial modulation of the inter-species scattering length in the binary Bose-Einstein condensates, in the form of a two-dimensional axisymmetric Gaussian, shaped by means of the optical Feshbach resonance. The corresponding effective nonlinear trapping potential supports completely stable elongated QDs with vorticity $S=0$ and partly stable families of elongated QDs with $S=1,2,3,4$ (other nonlinear systems do not maintain stability of vortex droplets with $\geq 2$). We systematically analyze effects of the amplitude and width of the Gaussian modulation, as well as the total number of atoms, on the shape and stability of the QDs, some effects being explained analytically. Collisions between identical QDs with $S=1$ moving in opposite directions along the central axis leads to their merger into still more elongated breathing QDs with the same vorticity, while collisions between QDs with $S=\pm 1$ are quasi-elastic. Moving modulation profiles are able to adiabatically rotate the trapped elongated QDs. Application of a torque to the vector QD sets in the gyroscopic regime of robust precession, which realizes a macroscopic spin-orbit-coupling effect.
I Introduction
@@
Vortices with an elongated three-dimensional (3D) structure have drawn much interest in astrophysics McWilliams1994a ; astro and geophysics McWilliams1994b , plasmas plasma1 ; plasma2 ; plasma3 , superconductors Dolan1989 ; Odobesko2020 ; Hung2012 ; Olszewski2018 , optics Freund1997 ; Sharma2011 , and Bose-Einstein condensates (BECs) Stockhofe2011 ; Liu2013 ; McEndoo2010 ; OKtel2004 ; Engels2002 ; Tikhonenkov2008 ; Tan2023 . However, the realization of elongated self-localized vortices - those stabilized solely by intrinsic interactions - has remained elusive. Quantum droplets (QDs) LHY ; Petrov2015 ; Petrov2016 ; Luca1 ; Jrgensen2018 ; Cabrera2018 ; Ferrier-Barbut2018 ; Cheiney2018 ; Inguscio2018 ; Luca2 ; Yin2021 ; dlw2021 ; Luo2021 ; Cui2025 offer a promising platform to address this challenge, as they inherently support the formation of self-localized vortices without external linear confinement, stable 2D and 3D QDs with embedded vorticity and more complex vortex patterns have been predicted Kartashov2018 ; QDvortex2D ; Tengstrand2019 ; Chen2024 ; Li2024FOPview .
The most intuitive solution lies in dipolar QDs where the stabilization emerges from the competitive balance between anisotropic dipole-dipole interactions (DDIs) Stuhler2005 ; Lahaye2009 and beyond mean-field effects. Although such systems exhibit intrinsic anisotropy Ferrier-Barbut2016 ; Schmitt2016 ; Santos ; Wang2020 ; Bottcher2021 ; Yin2022 , elongated vortex structures have proven fundamentally unstable Cidrim2018 , permitting only short-axis-aligned vortex configurations which the vorticity axis is perpendicular to the direction of the dipolar polarization Li2024PRL ; Li2024FOP .
An alternative approach within quantum droplets emerges from two-component atomic systems, which can sustain stable vortices but their inherent isotropic interactions prevent spontaneous elongation Kartashov2018 . This isotropic interaction is partly due to the dual-component stabilization mechanism, which utilizes magnetic Feshbach resonance (MFR) with dc magnetic field which alters the s-wave scattering length $a_{s}$ of the inter-atomic interaction to achieve uniform modulation of the interaction Tiesinga1993 ; Moerdijk1995 ; Courteille1998 ; Bertozzi1998 ; Inouye1998 ; Timmermans1999 ; Duine2004 ; kohler2006 ; Jones2006 ; Randy ; Chin2010 . In stark contrast, the characteristic feature of optical Feshbach resonance (OFR) - its ability to spatially inhomogeneously modulate scattering lengths Fisher1989 ; Fedichev1996 ; Caputo ; Saito2003 ; Qi2011 ; Chien2012 - reveals an extraordinary possibility for engineering elongated vortex states. It is worth noting that recent advancements in optical manipulation demonstrate OFR’s capacity for micro scale spatial density modulation in quantum droplets Schafer ; Fu , enabling precise morphology design, tailored macroscopic mechanical properties, and novel approaches for droplet dynamics manipulation. This technological progress establishes OFR as a transformative tool for quantum droplet engineering, particularly in realizing elongated vortex configurations.
The objective of the present work is to predict stable vortex QDs with a “slender” (elongated) shape in the binary condensate with the contact interactions, using an effective nonlinear potential RMP (alias pseudopotential pseudo ), that may be induced by a spatially modulated interaction strength. In turn, the modulation can be imposed by means of the spatially inhomogeneous Feshbach resonance - OFR, it can be readily induced by spatially non-uniform optical fields with inhomogeneous intensity $I\left(x,y,z\right)$, which can be used to shape effective nonlinear potentials through the local OFR relation,
@@
where $a_{\mathrm{opt}}$ is the OFR-induced change of the scattering length, $\left(x,y,z\right)$ are spatial coordinates, and sign minus implies that the resonant laser illumination induces the inter-atomic attraction force, which can be achieved by adjusting the detuning of the laser Theis2004 ; Bauer2009 ; Yamazaki2010 ; RMP ; Yan2013 .
II Model
@@
where $L$ is the beam waist, and $-a^{(0)}<0$ is the amplitude of the OFR-induced modulation of the local scattering length. The objective is to use the axisymmetric profile (2) to stabilize vortex solitons with the angular momentum directed along the $z$ axis and a “slender” (elongated) shape. The setting will be also used to study more sophisticated dynamical regimes, such as gyroscopic motion (precession) of the vortex soliton.
Our study, referring to the use of 39K atoms, implies large spectral detuning between the OFR wavelength and that corresponding to the atomic dipole transition, which makes it possible to neglect a linear potential induced by the OFR field Banerjee ; Campbell . In this case, $\psi_{1,2}$ are the wave functions of the two species, the total energy of the binary BEC being:
$E=\int_{V}\left[\sum_{j}\left(\frac{\hbar^{2}|\nabla\psi_{j}|^{2}}{2m}+\frac{\mathcal{G}_{jj}}{2}|\psi_{j}|^{4}\right)+\mathcal{G}_{12}|\psi_{1}|^{2}|\psi_{2}|^{2}\right.$
$\phantom{=\;\;}\left.+\frac{8m^{3/2}}{15\pi^{2}\hbar^{3}}\left(\sum_{j}{\mathcal{G}_{jj}}|\psi_{j}|^{2}\right)^{5/2}\right]\ ,\ \ j=1,2.$ (3)
Here $\hbar$ and $m$ are the Planck’s constant and atomic mass, while $\mathcal{G}_{jj}=4\pi\hbar^{2}a_{jj}/{m}$ ($j=1,2$) and $\mathcal{G}_{12}=4\pi\hbar^{2}a_{12}/{m}$ are the strengths of the intra- and inter-component mean-field interactions, respectively. Here, $a_{11},a_{22}$ and $a_{12}$ are the corresponding $s$-wave scattering lengths of atomic collisions, and the last term of Eq. (3) accounts for the LHY correction Petrov2015 ; Fischer2006 ; Inguscio2018 .
For this system, the minimization of mean-field energy is attained at $n_{1}/n_{2}=\sqrt{a_{22}/a_{11}}$ (where $n_{1,2}=|\psi_{1,2}|^{2}$) Petrov2015 ; Inguscio2018 . Here, we consider the symmetric case with $a_{22}=a_{11}$, which leads to the condition of $n_{2}=n_{1}$, hence the two wave-function coalesce into a single one as $\psi_{1}=\psi_{2}\equiv\psi/\sqrt{2}$. We consider the setup with the inter-component scattering length engineered by means of OFR so as to set $a_{12}=a_{\mathrm{bg}}+a_{\mathrm{opt}}$, where $a_{\mathrm{bg}}<0$ is the background value of the scattering length in the absence of the laser illumination. It is adjusted by the additional magnetic Feshbach resonance as $a_{\mathrm{bg}}=-a_{11}$, to cancel the spatially uniform scattering length (the detailed consideration of the two-component asymmetry is presented in the Appendix A). Thus, energy (3) is simplified to
$E=\int_{V}\left(\frac{\hbar^{2}}{2m}|\nabla\psi|^{2}+\frac{\pi\hbar^{2}a_{\mathrm{opt}}}{m}|\psi|^{4}+\frac{2}{5}\Gamma|\psi|^{5}\right),$ (4)
where $\Gamma=4m^{3/2}\mathcal{G}_{11}^{5/2}/3\pi^{2}\hbar^{3}$. Accordingly, the dynamics is governed by the single Gross-Pitaevskii (GP) equation:
@@
with angular coordinate $\theta$ and positive real function $\phi\left(\rho,z\right)$ satisfying the stationary equation
@@
III Characteristics of the Elongated states
To quantify the slenderness of 3D vortex QDs produced by Eq. (5), we define the aspect ratio, $\mathcal{A}=D_{\mathrm{v}}/D_{\mathrm{h}}$, where $D_{\mathrm{v}}$ and $D_{\mathrm{h}}$ are, respectively, vertical and horizontal lengths of the QDs. First, the numerical solution of the stationary version of Eq. (5) for fundamental QDs [with $S=0$ in Eq. (6)] demonstrates, in Figs. 1(a1-a3), the effect of the nonlinear trapping potential on the QD’s shape. In Fig. 1(a1), the aspect ratio is $\mathcal{A}=1$ for the isotropic fundamental QD in the free space ($L=\infty$), and $(a^{(0)},N)=(20a_{0},10^{6})$. As shown in Figs. 1(b1,b2), the aspect ratio increases (i.e., the fundamental QD elongates) with the decrease of radius $L$ of the holding beam in Eq. (2) and increase of $N$. The numerical solution of the time-dependent equation (5) (not shown here in detail) demonstrates that the entire family of the elongated QDs with $S=0$ is stable. Note, in particular, that the $\mu(N)$ dependence plotted in Fig. 1(b2) satisfies the well-known necessary stability condition, $d\mu/dN<0$ (the Vakhitov-Kolokolov criterion Vakhitov1973 ; Berge ).
@@ III Characteristics of the Elongated states
IV Dynamics
The independence of the modulation profile (2) on $z$ makes it possible to generate QDs freely moving along the $z$ axis, and thus simulate Eq. (5) for collisions between the QDs moving in opposite directions, initiated by the input, $\psi_{0}\left(\rho,z\right)=\psi(\rho,z-z_{d})\mathrm{e}^{-i\eta z}+\psi(\rho,z+z_{d})\mathrm{e}^{i\eta z}$, where $2z_{d}$ is the initial separation between the QDs, and ±$\pm\eta$ are the opposite kicks which set the droplets in motion. Figures 3(a1-a4) show that the collision between the QDs with identical vorticities $S=1$ is inelastic, resulting in the merger of the vortex QDs into a single oscillatory one (breather). The breathing dynamics is robust and quasi-harmonic, lasting, at least, for $>100\ \mathrm{ms}$, as is illustrated in Fig. 3(b) a sequence of $5$ cycles within the time interval of $50\thicksim 75\ \mathrm{ms}$. The breathing frequency is nearly independent of the collision velocity, suggesting that it is an eigenfrequency of a QD’s intrinsic mode. The collision leads to full destruction of the vortex QDs if the initial kick exceeds a certain critical value, which is $\eta_{\mathrm{cr}}=1.2\ \upmu\mathrm{m^{-1}}$. Collisions between zero-vorticity QDs produce similar results (not shown here in detail). On the other hand, for the QDs with opposite vorticities, $S=\pm 1$, the simulations demonstrate that the collision is quasi-elastic, leading to temporary fusion into a single droplet, which then splits back into original moving vortex QDs. The latter result is explained by the fact that the opposite vorticities attenuate the coherent interaction between the colliding objects. It is shown in Fig. 3(c) that the holding beam rotating around the $y$-axis in the $\left(xz\right)$-plane at angular velocity $\omega=0.163\ \mathrm{rad/ms}$, which means the replacement of the trapping profile (2) by
@@ IV Dynamics
Figure 4: (a-c) Typical examples of the vortex droplets and their phases diagrams (at $z=0$) with $(L,a^{(0)},N)=(8\ \upmu\mathrm{m},20a_{0},2\times 10^{6})$. (d) The chemical potential $\mu$ of the QDs with $S=0\sim 4$ vs. $N$, with solid and dashed lines representing stable and unstable solutions, respectively.
V Higher-order vortex states
@@
VI Conclusion
We have proposed to use the OFR, modulating the inter-component atomic $s$-wave scattering length in the binary BEC, to induce the effective 2D nonlinear trapping potential with the Gaussian profile. It supports elongated QDs, including ones with embedded integer vorticity $S$, which are stabilized by the GP-LHY equations. We have systematically explored the effects of the modulation amplitude and width, as well as the total number of atoms, on the shape and stability of the QDs. The fundamental QDs with $S=0$ are completely stable, while families of the vortex QDs with $S=1,2,3,$ and $4$ are partly stable. Collisions of identical elongated vortex QDs with $S=1$ lead to their merger into breathers with the same vorticity, $S=1$, while collisions between the elongated QDs with opposite vorticities, $S=\pm 1$, are quasi-elastic. The moving OFR modulation profile can rotate the elongated QDs in the abiabatic regime. The application of the torque to the trapped elongated vortex QD sets it in a state of stable gyroscopic motion (precession).
As an extension of the analysis, it may be interesting to consider more complex holding potentials, such as anisotropic ones, as well as those combining the nonlinear trap and vortex laser beams. Such modes with long inner voids can be used, in the context of the currently developing field matter-wave atomtronics, as conduits (waveguides) for linear atomic jets, similar to the concept of ”light guided by light”, which is wellknown in nonlinear optics Demircan ; Vasa ; Fang . Magnetic-gradient/radiofrequency-mediated scattering-length control, combined with droplet-creation schemes, could simplify future experiments Carli ; Sanz ; Lavoine .